# On $\nu$ faces of partial alternating sign matrix polytopes

### v1

@@ 1. Introduction and background
The ASM-CRY polytope is an alternating sign matrix analogue of the CRY polytope: a face of the Birkhoff polytope studied by Chan, Robbins, and Yuen [3] whose volume is given as a product of Catalan numbers [19]. It was shown in [9] that the ASM-CRY polytope is integrally equivalent to an order polytope, which provided a proof of its lovely combinatorial properties, such as counting formulas for its volume and vertices. More generally, [9] showed that order polytopes of strongly planar posets are flow polytopes. Subsequent work [2, 8, 7, 5, 17] built upon this connection between order and flow polytopes, but there was no interpretation in terms of alternating sign matrices. In this paper, we reunite these topics. In particular, we were inspired by the polytope perspective on the $\nu$-Tamari lattice in [17] to define and study an analogue of the ASM-CRY polytope of [9] using the partial alternating sign matrix polytope of [4]. Furthermore, we show this polytope is both an order and flow polytope.
@@
Alternating sign matrices arose in the study of the $\lambda$-determinant [13]. They have a lovely enumeration formula [10, 18, 6] and connections to statistical physics [6]. The alternating sign matrix polytope was defined as above by Striker in [16], where she also found an inequality description, enumerated the facets, and studied its projection to the permutohedron. Independently, Behrend and Knight defined this polytope in terms of inequalities and proved the vertex description [1]. In addition, they studied lattice points in the $r$th-dilate of the alternating sign matrix polytope, which they called higher spin alternating sign matrices.
In [9], a face of the alternating sign matrix polytope with some prescribed zeros in certain parts of the matrix was studied and shown to be integrally equivalent to an order polytope. We give the definition of this polytope and state the relevant theorems below. (Our convention differs from that of [9] by vertical reflection of the matrix, to match with the conventions of this paper.)
@@
The normalized volume of $\mathrm{ASMCRY}(\lambda,n)$ is given by the number of linear extentions $e(P(\delta_{n}/\lambda))$. Its Ehrhart polynomial $L_{\mathrm{ASMCRY}(\lambda,n)}(t)$ equals the order polynomial $\Omega(P(\delta_{n}/\lambda),t+1)$.
Moreover, when $\lambda=\emptyset$, $\mathrm{ASMCRY}(n)$ has $\mathrm{Cat}(n)=\frac{1}{n+1}\binom{2n}{n}$ vertices, its normalized volume is given by the number of standard Young tableaux of shape $\delta_{n}$, and its Ehrhart polynomial is $L_{\mathrm{ASMCRY}(n)}(t)=\Omega(P(\delta_{n}),t+1)=\prod_{1\leq i<j\leq n}\frac{2t+i+j-1}{i+j-1}$.
The partial alternating sign matrix polytope $\mathrm{PASM}(m,n)$ was introduced and studied in [4]. This polytope is defined as the convex hull of partial alternating sign matrices, a natural extension of alternating sign matrices, whose definition we give in Definition 2.1. The authors proved an inequality desription of the partial alternating sign matrix polytope. They also enumerated the facets and characterized the face lattice, in an analogous way to that of the alternating sign matrix polytope.
In this paper, we study polytopes $\mathrm{PASM}(\nu/\lambda,m,n)$ (see Definition 2.6) contained in partial alternating sign matrix polytopes, with certain entries set equal to 00, analogous to the ASM-CRY polytope of Definition 1.2. We find similar connections to order and flow polytopes and the results of [9] discussed above. Our theorems are not straightforward generalizations of results of [9], though. We show in Proposition 5.11 that when restricted to square matrices and the appropriate partition shapes, $\mathrm{PASM}(\delta_{n}/\lambda,n,n)$ is not equal to $\mathrm{ASM}(\lambda,n)$, but the polytopes are integrally equivalent.
@@
Let $\lambda,\mu,\nu$ be partitions inside the partition $n^{m}$ such that $\lambda\subseteq\mu\subseteq\nu$. Then $\mathrm{PASM}(\nu/\lambda,m,n)$ is given by the inequalities:
@@
Our final main result, in Section 5, shows that $\mathrm{PASM}(\nu/\lambda)$ is integrally equivalent to an order polytope (and thus a flow polytope, see Corollary 5.10).
@@
The $(\nu/\lambda)$-partial ASM polytope $\mathrm{PASM}(\nu/\lambda)$ is integrally equivalent to the order polytope $\mathcal{O}(P(\nu/\lambda))$.
This leads to the following result on the volume and Ehrhart polynomial of $\mathrm{PASM}(\nu/\lambda)$.
@@
The normalized volume of $\mathrm{PASM}(\nu/\lambda)$ is $e(P(\nu/\lambda))$ and its Ehrhart polynomial is $L_{\mathrm{PASM}(\nu/\lambda)}(t)=\Omega(P(\nu/\lambda),t+1)$. In particular, $e(P(\nu/\lambda))$ is the number of skew standard Young tableaux, enumerated by the Naruse hook-length formula
@@ 2. Definition of the $(\nu/\lambda)$-partial ASM polytope
We now define a set of matrices the associated polytope that will be the main focus of study in this paper.
@@ 3. Inequality description
In this section, we define the type of partitions being used and prove an inequality description for the polytope $\mathrm{PASM}(\nu/\lambda,m,n)$.
@@
Let $\lambda,\mu,\nu$ be partitions inside the partition $(n-1)^{m-1}$ such that $\lambda\subseteq\mu\subseteq\nu$. Then $\mathrm{PASM}(\nu/\lambda,m,n)$ is given by the inequalities:
@@
    $\sum_{1\leq j\leq n}X_{ij}=0\text{ for }2\leq i\leq m\text{ and }\sum_{1\leq i\leq m}X_{ij}=0\text{ for }2\leq j\leq n$ (3.4)
@@
Fix $m,n>0$ and partitions $\lambda\subseteq\nu\subset(n-1)^{m-1}$. Since $m$ and $n$ are fixed, we suppress them in the notation on individual matrices, for clarity.
@@
Since $\lambda\subseteq\mu$, and the $1$’s and $-1$’s in $M^{\mu}$ have been placed outside of $\mu$, all entries in $\lambda$ are equal to 00. So (3.5) is satisfied. Also, since $\mu\subseteq\nu$ and all $1$’s and $-1$’s in $M^{\mu}$ have been placed at most one box south and/or east of $\mu$, all entries southeast of the border strip associated to $\nu$ are equal to 00. So (3.6) is satisfied. ∎
@@ 4. The $(\nu/\lambda)$-partial ASM polytope is a face of the partial ASM polytope
Given an $m\times n$ matrix $X$, we define a labeled graph, $\widehat{X}$, which is a labeling of the vertices and edges of $\Gamma_{(m,n)}$. The horizontal edges from $(i,j)$ to $(i,j+1)$ are each labeled by the corresponding row partial sum $\sum_{j^{\prime}=1}^{j}X_{ij^{\prime}}$ ($1\leq i\leq m,1\leq j\leq n$). Similarly, the vertical edges from $(i,j)$ to $(i+1,j)$ are each labeled by the corresponding column partial sum $\sum_{i^{\prime}=1}^{i}X_{i^{\prime}j}$ ($1\leq i\leq m,1\leq j\leq n$).
@@
Given $M\in\mathrm{PASM}_{m,n}$, the basic sum-labeling $g(M)$ of ${\Gamma}_{(m,n)}$ is the labeling of the edges of $\Gamma_{(m,n)}$ with 00 or $1$ such that the edge labels equal the corresponding edge labels of $\widehat{M}$. Given a collection of partial alternating sign matrices =$\mathcal{M}=\{M_{1},M_{2},\dots,M_{r}\}\subseteq\mathrm{PASM}_{m,n}$, define the sum-labeling $g(\mathcal{M})=\displaystyle\bigcup_{i=1}^{r}g(M_{i})$, where the union is defined as the labeling of $\Gamma_{(m,n)}$ whose edge labels are given by the union of the corresponding edge labels of the $g(M_{i})$. The set of sum-labelings of $\Gamma_{(m,n)}$ is the set of $g(\mathcal{M})$ over all collections =$\mathcal{M}=\{M_{1},M_{2},\dots,M_{r}\}\subseteq\mathrm{PASM}_{m,n}$, $(0\leq r)$.
@@
Figure 3. Examples of partial sum labelings, corresponding to the matrices in Figure 2, where $\mu$ is between $\lambda=(3,1)$ and $\nu=(4,2,2)$.
@@
Using the characterization of faces of $\mathrm{PASM}(m,n)$ as sum labelings in the above theorem, we show that $\mathrm{PASM}(\nu/\lambda,m,n)$ is a face of $\mathrm{PASM}(m,n)$.
@@
By Theorem 4.4, the faces of $\mathrm{PASM}(m,n)$ are in bijection with sum labelings of $\Gamma_{(m,n)}$. Since, by Definition 2.6, $\mathrm{PASM}(\nu/\lambda,m,n)$ is the convex hull of the partial alternating sign matrices $\mathrm{PASM}^{\nu/\lambda}_{m,n}$, we need only show there exists a sum labeling whose contained basic sum labelings correspond exactly to these partial alternating sign matrices. We give this sum labeling explicitly.
@@
To create the desired sum labeling, label the edges in $\left(\bigcup_{\lambda\subseteq\mu\subseteq\nu}E_{\mu}\right)-(E_{\lambda}\cap E_{\nu})$ with $\{0,1\}$ (the blue lines in Figure 4). Label the edges in the intersection $E_{\lambda}\cap E_{\nu}$ with $1$. Label all remaining edges of $\Gamma_{(m,n)}$ with 00. See Example 4.6 for an example of an explicit construction of $E_{\lambda}$ and $E_{\nu}$.
@@
By construction, this sum labeling has $|\nu|-|\lambda|$ regions (exactly the boxes which correspond to $\nu/\lambda$ in $\Gamma_{(m,n)}$). Thus by Theorem 4.4, the dimension of $\mathrm{PASM}(\nu/\lambda,m,n)$ equals $|\nu|-|\lambda|$. ∎
@@
In Figure 4, we have constructed the basic sum labeling corresponding to $\mathrm{PASM}_{\nu/\lambda}$ where $\lambda=(3,1)$ and $\nu=(4,2,2)$. We show this construction as follows; see Figure 5.
@@
Figure 4. The union of the sum labelings in Figure 2, with $\mu$ between $\lambda=(3,1)$ and $\nu=(4,2,2)$.

### v2

@@ 1. Introduction and background
The ASM-CRY polytope is an alternating sign matrix analogue of the CRY polytope: a face of the Birkhoff polytope studied by Chan, Robbins, and Yuen [3] whose volume is given as a product of Catalan numbers [19]. It was shown in [9] that the ASM-CRY polytope is integrally equivalent to an order polytope, which provided a proof of its lovely combinatorial properties, such as counting formulas for its volume and vertices. More generally, [9] showed that order polytopes of strongly planar posets are flow polytopes. Subsequent work [2, 8, 7, 5, 17] built upon this connection between order and flow polytopes, but there was no interpretation in terms of alternating sign matrices. In this paper, we reunite these topics. In particular, we were inspired by the polytope perspective on the $\nu$-Tamari lattice in [17] to define and study an analogue of the ASM-CRY polytope of [9] using the partial alternating sign matrix polytope of [4]. Furthermore, we show this polytope is both an order polytope and a flow polytope.
@@
Alternating sign matrices arose in the study of the lambda-determinant [13]. They have a nice enumeration formula [10, 18, 6] and connections to statistical physics [6]. The alternating sign matrix polytope was defined as above by Striker in [16], where she also found an inequality description, enumerated the facets, and studied its projection to the permutohedron. Independently, Behrend and Knight defined this polytope in terms of inequalities and proved the vertex description [1]. In addition, they studied lattice points in the $r$th-dilate of the alternating sign matrix polytope, which they called higher spin alternating sign matrices.
In [9], a face of the alternating sign matrix polytope with some prescribed zeros in certain parts of the matrix was studied and shown to be integrally equivalent to an order polytope. We give the definition of this polytope and state the relevant theorems below. (Note our convention differs from that of [9] by vertical reflection of the matrix.)
@@
The main results about this polytope from [9] are the following; see Section 5 for definitions of the notation.
@@
The normalized volume of $\mathrm{ASMCRY}(\lambda,n)$ is given by the number of linear extensions $e(P(\delta_{n}/\lambda))$. Its Ehrhart polynomial $L_{\mathrm{ASMCRY}(\lambda,n)}(t)$ equals the order polynomial $\Omega(P(\delta_{n}/\lambda),t+1)$.
Moreover, $\mathrm{ASMCRY}(\emptyset,n)$ has $\mathrm{Cat}(n)=\frac{1}{n+1}\binom{2n}{n}$ vertices, its normalized volume is given by the number of standard Young tableaux of shape $\delta_{n}$, and its Ehrhart polynomial is $L_{\mathrm{ASMCRY}(\emptyset,n)}(t)=\Omega(P(\delta_{n}),t+1)=\prod_{1\leq i<j\leq n}\frac{2t+i+j-1}{i+j-1}$.
The partial alternating sign matrix polytope $\mathrm{PASM}(m,n)$ was introduced and studied in [4]. This polytope is defined as the convex hull of partial alternating sign matrices, a natural extension of alternating sign matrices, whose definition we give in Definition 2.1. The authors of [4] proved an inequality description of the partial alternating sign matrix polytope. They also enumerated the facets and characterized the face lattice, in an analogous way to that of the alternating sign matrix polytope.
In this paper, we study polytopes $\mathrm{PASM}(\nu/\lambda,m,n)$ (see Definition 2.6) contained in partial alternating sign matrix polytopes, with certain entries set equal to 00, analogous to the ASM-CRY polytope of Definition 1.2. We find similar connections to order and flow polytopes and the results of [9] discussed above. Our theorems are not straightforward generalizations of results of either [4] or [9]. For instance, our proof of Theorem 3.2 involves a novel induction. Also, we show in Proposition 5.11 that when restricted to square matrices and the appropriate partition shapes, $\mathrm{PASM}(\delta_{n}/\lambda,n,n)$ is not equal to $\mathrm{ASMCRY}(\lambda,n)$, but the polytopes are integrally equivalent.
@@
Let $\lambda,\nu$ be partitions inside the partition $n^{m}$ such that $\lambda\subseteq\nu$. Then $\mathrm{PASM}(\nu/\lambda,m,n)$ is given by the inequalities:
@@
Our final main result, in Section 5, shows that $\mathrm{PASM}(\nu/\lambda,m,n)$ is integrally equivalent to an order polytope of a strongly planar graph (and thus a flow polytope, see Corollary 5.10).
@@
The $(\nu/\lambda)$-partial ASM polytope $\mathrm{PASM}(\nu/\lambda,m,n)$ is integrally equivalent to the order polytope $\mathcal{O}(P(\nu/\lambda))$.
This leads to the following result on the volume and Ehrhart polynomial of $\mathrm{PASM}(\nu/\lambda,m,n)$. Note for a poset $P$, $\Omega(P,t)$ denotes the number of order-preserving maps $\eta:P\rightarrow\{1,\ldots,t\}$.
@@
The normalized volume of $\mathrm{PASM}(\nu/\lambda,m,n)$ is $e(P(\nu/\lambda))$ and its Ehrhart polynomial is $L_{\mathrm{PASM}(\nu/\lambda,m,n)}(t)=\Omega(P(\nu/\lambda),t+1)$. In particular, $e(P(\nu/\lambda))$ is the number of skew standard Young tableaux, enumerated by the Naruse hook-length formula
@@ 2. Definition of the $(\nu/\lambda)$-partial ASM polytope
We now define the set of matrices of the associated polytope that will be the main focus of study in this paper.
@@ 3. Inequality description
In this section, we prove an inequality description for the polytope $\mathrm{PASM}(\nu/\lambda,m,n)$.
@@
Let $\lambda,\nu$ be partitions inside the partition $(n-1)^{m-1}$ such that $\lambda\subseteq\nu$. Then $\mathrm{PASM}(\nu/\lambda,m,n)$ is given by the inequalities:
@@
    $\sum_{j=1}^{n}X_{ij}=0\text{ for }2\leq i\leq m\text{ and }\sum_{i=1}^{m}X_{ij}=0\text{ for }2\leq j\leq n$ (3.4)
@@
Fix $m,n>0$ and partitions $\lambda\subseteq\nu\subset(n-1)^{m-1}$.
@@
Since $\lambda\subseteq\mu$, and the $1$’s and $-1$’s in $M^{\mu}$ have been placed outside of $\mu$, all entries in $\lambda$ are equal to 00. So (3.5) is satisfied. Also, since $\mu\subseteq\nu$ and all $1$’s and $-1$’s in $M^{\mu}$ have been placed at most one box south and/or east of $\mu$, all entries southeast of the border strip associated to $\nu$ are equal to 00. So (3.6) is satisfied.
Now we wish to show that $\mathcal{P}(\nu/\lambda,m,n)$ is contained in $\mathrm{PASM}(\nu/\lambda,m,n)$. Suppose that $X\in\mathcal{P}(\nu/\lambda,m,n)$. By (3.1) and (3.2), $X$ is also in $\mathrm{PASM}(m,n)$. So we can write it as the convex combination of some partial ASMs: $X=\displaystyle\sum_{k=1}^{w}c_{k}M^{k}$ with $c_{k}>0$ and $\sum_{k=1}^{w}c_{k}=1$.
Suppose one of these partial alternating sign matrices does not satisfy the first equality in (3.3). Without loss of generality, say it is $M^{1}$. Since it is a partial alternating sign matrix that does not have a first column total sum of 1, then it must have a first column sum of 0. So
$\sum_{i=1}^{m}X_{i1}=\displaystyle\sum_{i=1}^{m}\displaystyle\sum_{k=1}^{w}c_{k}M^{k}_{i1}=\displaystyle\sum_{k=1}^{w}c_{k}\displaystyle\sum_{i=1}^{m}M^{k}_{i1}=c_{1}\displaystyle\sum_{i=1}^{m}M^{1}_{i1}+\displaystyle\sum_{k=2}^{w}c_{k}\displaystyle\sum_{i=1}^{m}M^{k}_{i1}=c_{1}\cdot 0+\displaystyle\sum_{k=2}^{w}c_{k}\cdot 1<1.$
This is a contradiction. We can do a similar proof for the second equality in (3.3). So each $M^{k}$ satisfies (3.3).
Now suppose one of these partial alternating sign matrices does not satisfy the first equality in (3.4) for a fixed value of $i$ with $2\leq i\leq m$. Without loss of generality, say this matrix is $M^{1}$. Since this is a partial alternating sign matrix that does not have an $i$th row total sum of 0, then it must have an $i$th row total sum of 1. So
$\sum_{j=1}^{n}X_{ij}=\displaystyle\sum_{j=1}^{n}\displaystyle\sum_{k=1}^{w}c_{k}M^{k}_{ij}=\displaystyle\sum_{k=1}^{w}c_{k}\displaystyle\sum_{j=1}^{n}M^{k}_{ij}=c_{1}\displaystyle\sum_{j=1}^{n}M^{1}_{ij}+\displaystyle\sum_{k=2}^{w}c_{k}\displaystyle\sum_{j=1}^{n}M^{k}_{ij}=c_{1}\cdot 1+\displaystyle\sum_{k=2}^{w}c_{k}\cdot 0=c_{1}>0.$
This is a contradiction. We can do a similar proof for the second equality in (3.4). So each $M^{k}$ satisfies (3.4).
Now suppose one of these partial alternating sign matrices, call it $M^{1}$, does not satisfy the (3.5) for some $j<\lambda_{i}$. Suppose this is the smallest such value of $j$ with this $i$. So this is a partial ASM with $i$th row partial sum $\sum_{j^{\prime}=1}^{j}M^{1}_{ij}=1$ but $\sum_{j^{\prime}=1}^{j-1}M^{k}_{ij}=0$ for all $k$. But since $X$ satisfies (3.5), $\sum_{j^{\prime}=1}^{j}X_{ij}=0$. So
$\sum_{j^{\prime}=1}^{j}X_{ij}=\displaystyle\sum_{j^{\prime}=1}^{j}\displaystyle\sum_{k=1}^{w}c_{k}M^{k}_{ij}=\displaystyle\sum_{k=1}^{w}c_{k}\displaystyle\sum_{j^{\prime}=1}^{j}M^{k}_{ij}=c_{1}\displaystyle\sum_{j^{\prime}=1}^{j}M^{1}_{ij}+\displaystyle\sum_{k=2}^{w}c_{k}\displaystyle\sum_{j^{\prime}=1}^{j}M^{k}_{ij}=c_{1}\cdot 1+\displaystyle\sum_{k=2}^{w}c_{k}\cdot 0=c_{1}>0.$
This is a contradiction. So each $M^{k}$ satisfies (3.5). We can do a similar proof for (3.6) by choosing the largest such $j$ rather than the smallest.
Thus we have shown both $\mathcal{P}(\nu/\lambda,m,n)\subseteq\mathrm{PASM}(\nu/\lambda,m,n)$ and $\mathrm{PASM}(\nu/\lambda,m,n)\subseteq\mathcal{P}(\nu/\lambda,m,n)$, verifying these polytopes are equal. ∎
@@ 4. The $(\nu/\lambda)$-partial ASM polytope is a face of the partial ASM polytope
Given an $m\times n$ matrix $X$, we define a labeled graph, $g(X)$, which is a labeling of the edges of $\Gamma_{(m,n)}$. The horizontal edges from $(i,j)$ to $(i,j+1)$ are each labeled by the corresponding row partial sum $\sum_{j^{\prime}=1}^{j}X_{ij^{\prime}}$ ($1\leq i\leq m,1\leq j\leq n$). Similarly, the vertical edges from $(i,j)$ to $(i+1,j)$ are each labeled by the corresponding column partial sum $\sum_{i^{\prime}=1}^{i}X_{i^{\prime}j}$ ($1\leq i\leq m,1\leq j\leq n$).
@@
Given $M\in\mathrm{PASM}_{m,n}$, we call $g(M)$ the basic sum-labeling of $M$. Given a collection of partial alternating sign matrices =$\mathcal{M}=\{M_{1},M_{2},\dots,M_{r}\}\subseteq\mathrm{PASM}_{m,n}$, define the sum-labeling $g(\mathcal{M})=\displaystyle\bigcup_{i=1}^{r}g(M_{i})$, where the union is defined as the labeling of $\Gamma_{(m,n)}$ whose edge labels are given by the union of the corresponding edge labels of the $g(M_{i})$. The set of sum-labelings of $\Gamma_{(m,n)}$ is the set of $g(\mathcal{M})$ over all collections =$\mathcal{M}=\{M_{1},M_{2},\dots,M_{r}\}\subseteq\mathrm{PASM}_{m,n}$, $(0\leq r)$.
@@
Figure 3. Examples of basic sum-labelings, corresponding to the matrices in Figure 2, where $\mu$ is between $\lambda=(3,1)$ and $\nu=(4,2,2)$.
@@
Using the characterization of faces of $\mathrm{PASM}(m,n)$ as sum-labelings in the above theorem, we show that $\mathrm{PASM}(\nu/\lambda,m,n)$ is a face of $\mathrm{PASM}(m,n)$.
@@
By Theorem 4.4, the faces of $\mathrm{PASM}(m,n)$ are in bijection with sum-labelings of $\Gamma_{(m,n)}$. Since, by Definition 2.6, $\mathrm{PASM}(\nu/\lambda,m,n)$ is the convex hull of the partial alternating sign matrices $\mathrm{PASM}^{\nu/\lambda}_{m,n}$, we need only show there exists a sum-labeling whose contained basic sum-labelings correspond exactly to these partial alternating sign matrices. We give this sum-labeling explicitly.
@@
To create the desired sum-labeling, label the edges in $\left(\bigcup_{\lambda\subseteq\mu\subseteq\nu}E_{\mu}\right)-(E_{\lambda}\cap E_{\nu})$ with $\{0,1\}$ (the blue lines in Figure 4). Label the edges in the intersection $E_{\lambda}\cap E_{\nu}$ with $1$. Label all remaining edges of $\Gamma_{(m,n)}$ with 00. See Example 4.6 for an example of an explicit construction of $E_{\lambda}$ and $E_{\nu}$.
@@
By construction, this sum-labeling has $|\nu|-|\lambda|$ regions (exactly the boxes which correspond to $\nu/\lambda$ in $\Gamma_{(m,n)}$). Thus by Theorem 4.4, the dimension of $\mathrm{PASM}(\nu/\lambda,m,n)$ equals $|\nu|-|\lambda|$. ∎
@@
In Figure 4, we have constructed the basic sum-labeling corresponding to $\mathrm{PASM}_{\nu/\lambda}$ where $\lambda=(3,1)$ and $\nu=(4,2,2)$. We show this construction as follows; see Figure 5.
@@
Figure 4. The union of the sum-labelings in Figure 2, with $\mu$ between $\lambda=(3,1)$ and $\nu=(4,2,2)$.